A sprinter is jogging at 4.0 m s-1 and accelerates at 10.0 m s-2 over 49.0 m. Find the velocity of the sprinter


1
Expert's answer
2021-02-23T10:00:06-0500

The distance traveled by the sprinter

"d=\frac{v_f^2-v_i^2}{2a}"

Hence, the final velosity of the sprinter

"v_f=\sqrt{v_i^2+2da}=\sqrt{4.0^2+2*49.0*10.0}=31.6\: \rm m/s"